OpenStudy (lizzie.tttt):

What is the slope of the graph of the equation below? 3x+y=5 A; 3 B; -3 C; 5 <-- My answer D; -5

OpenStudy (helpkay):

i use papa math trhis should be able to help yopu with your work http://www.mathpapa.com/algebra-calculator.html

OpenStudy (dayakar):

slope of a line ax+by+c=0 is slope (m) =- a/b

OpenStudy (dayakar):

can u do now

OpenStudy (lizzie.tttt):

Wait is it -3?

OpenStudy (dayakar):

3x+1y=5 here a= 3 and b=1 find m=

OpenStudy (lizzie.tttt):

3?

OpenStudy (dayakar):

why r u changing the answers m= -3

OpenStudy (helpkay):

wait what y=mx+b the b is your slope so it would be 5